When 0.344 mol of HCl is combined with enough water to make a 450.0 mL solution, the concentration of HCl is ________ M.

Question 8 options:


1) 

7.64 × 10-4


2) 

0.155


3) 

1.31


4) 

0.764


5) 

0.344


1
Expert's answer
2022-06-13T17:03:03-0400

CM — concentration of molarity = ?

n — amount of HCl = 0,344 mol

V — volume of solution = 450 ml = 0,45 l


CM = n / V = 0,344 / 0,45 = 0,764 mol/l


Answer:

4) 0,764